graph the equation 5x-y=5
Answer by checkley71(8403)

5X-Y=5 OR -Y=-5X+5 OR Y=5X-5 THUS A SLOPE OF 5 & A Y INTERCEPT OF -5
ANOTHER POINT CAN BE CALCULATED BY SETTING Y=0 OR 5X-0=5 OR X=5/5 OR X=1
NOW WE HAVE (0,-5) & (1,0) AS TWO POINTS ON THIS LINE.